With regard to achievement of stability of equilibrium in an isolated market assume that the market for string beans is found to have a lagged supply response such that the demand and supply function may be written as 2 Dt = aPt + b St = AP t-1 + B

Required: Derive the conditions for dynamic stability of equilibrium


1
Expert's answer
2021-09-29T18:01:31-0400

Given:

"Dt = aPt + b\\

St = APt-1 + B"

At the equilibrium level,

"Dt = St\\

aPt + b = APt-1 + B .....(1)"

If the equilibrium is stable it implies that the price level in all the periods is equal to the stable prices, i.e P*.

"=> aP^* + b = AP^* + B .....(2)"

Subtracting 1 from 2 we get,

"a (P^* - Pt) = A(P^* - Pt-1)\\

aP1 = AP2"

Where P1 = Deviation of Pt from P*

P2 = Deviation of Pt-1 from P*

"\frac{P_1}{P_2} = \frac{A}{a}" , where "\frac{A}{a}" is a constant denoted by C

"\frac{P_1}{P_2} =C"